i need help on these two questions
(3/2)x+.5x+7=0
3/2x+11/5x=1
i need to solve for the variables. thank you

Answer by Alan3354(69443)

i need help on these two questions
(3/2)x+.5x+7=0
Step 1, multiply thru by 2 to eliminate fractions.
================
3/2x+11/5x=1
If you mean
(3/2)x + (11/5)x = 1
Step 1, multiply thru by 10 to eliminate fractions.
You should use parentheses to eliminate ambiguities.
I suspect some tutors ignore problems that are ambiguous. I know I do.
================
i need to solve for the variables